Che volume V of the solid obtained by rotating the region bounded b y = 3 + sec(x), sxs y = 5; about y = 3

Find the volume of the solid
Answer #1

Y = 34 seccx) -I = XE Y5 3 Rotation about Y=3 4= 3+ secx) washer method using y=5 outer radius = R. (X) = 5-3 y=3 inner radiu